# STONE: Self-supervised Tonality Estimator

### v2

@@ 1 Introduction
A new large dataset of 5489 real-world music recordings, collected from the Free Music Archive (FMA) and annotated by an expert for 24 major and minor keys, is available for free download [^2] FMAK dataset: https://zenodo.org/records/12759100.
@@ 3.2 ChromaNet: a convnet with octave equivalence
We compose the convnet $f_{\bm{\theta}}$ with a non-trainable operator $g$ whose role is to guarantee octave equivalence. We roll the log-frequency axis into a spiral which makes a full turn at every octave, thus aligning coefficients in $f_{\bm{\theta}}(T_{c}\bm{x})$ of the form $(p\pm Qj)$ for integer $j$. This averaging process is a common operation in literature[26]. The operator $g$ sums these coefficients across octaves $j$ for each pitch class $q$ and applies a softmax transformation. Composing $f_{\bm{\theta}}$ and $g$ yields
$\bm{y}_{\bm{\theta}}[q]$ $=(g\circ f_{\bm{\theta}})(\bm{x})[q]$
$=\dfrac{\exp\left(\sum_{j=0}^{J-1}f_{\bm{\theta}}(\bm{x})[Qj+q]\right)}{\sum_{q^{\prime}=0}^{Q-1}\exp\left(\sum_{j=0}^{J-1}f_{\bm{\theta}}(\bm{x})[Qj+q^{\prime}]\right)},$ (1)
i.e., a vector with $Q$ nonnegative entries which sum to one. The equation above resembles the extraction of chroma features [27], hence the proposed name of ChromaNet.
@@ 3.3 DFT over key signature profiles
We propose the term “key signature profile” for the learned vector $\bm{y}_{\bm{\theta}}$. With $Q=12$, its discrete Fourier transform (DFT) is
@@ 3.4 Cross-power spectral density (CPSD)
Let us split the CQT matrix $\bm{x}$ into two disjoint time segments of equal length: $\bm{x}=(\bm{x}_{\mathrm{A}},\bm{x}_{\mathrm{B}})$. We denote the ChromaNet response for A by $\bm{y}_{\bm{\theta},{\mathrm{A}}}=(g_{Q}\circ f_{\bm{\theta}})(\bm{x}_{\mathrm{A}})$ idem for B. The circular cross-correlation between $\bm{y}_{\bm{\theta},{\mathrm{A}}}$ and $\bm{y}_{\bm{\theta},{\mathrm{B}}}$ is
@@ 4.1 Training on real-world unlabeled data
The self-supervised procedure above learns an approximately equivariant mapping from CQT to key signature profiles (KSP). After training, we observe informally that for each input $\bm{x}$, most of the softmax activation in the KSP $\bm{y}_{\bm{\theta}}$ is concentrated on a single pitch class. In other words, the nonconstrative loss $\mathcal{L}^{\mathrm{CPSD}}$ (Equation 9) is sparsity-promoting.
@@ 4.5 Results on key signature estimation
Feature engineering 1352 3438 31%
STONE $(\omega=7)$ 3587 931 74%
STONE $(\omega=1)$ 3883 751 78%
Supervised SOTA [21] 4090 755 81%
@@ 7 Conclusion
STONE learns key signature profiles (KSP) via noncontrastive self-supervised learning in the time–frequency domain. We have seen than a semi-supervised extension of STONE (semi-TONE) reduces expert annotation by 90% less at no loss of MIREX score compared to the fully supervised variant (sup-TONE). The primary limitation of our work resides in the inability of the STONE objective (CPSD, i.e., cross-power spectral density) to distinguish major keys from minor keys.re work will study how STONE can be adapted to other pitch-relative MIR tasks.

### v3

@@ 1 Introduction
A new large dataset of 5489 real-world music recordings, collected from the Free Music Archive (FMA) and annotated by an expert for 24 major and minor keys, is available for free download [^2] FMAK dataset (version 2): https://zenodo.org/records/12759100.
@@ 3.2 ChromaNet: a convnet with octave equivalence
We compose the convnet $f_{\bm{\theta}}$ with a non-trainable operator $g$ whose role is to guarantee octave equivalence. We roll the log-frequency axis into a spiral which makes a full turn at every octave, thus aligning coefficients in $f_{\bm{\theta}}(T_{c}\bm{x})$ of the form $(p\pm Qj)$ for integer $j$. The operator $g$ sums these coefficients across octaves $j$ for each pitch class $q$ and applies a softmax transformation. We obtain a $Q$-dimensional vector $\bm{y}_{\bm{\theta}}$ whose entries are nonnegative and sum to one. We propose to call this vector a “key signature profile” (KSP):
$\bm{y}_{\bm{\theta}}[q]$ $=(g\circ f_{\bm{\theta}})(T_{c}\bm{x})[q]$
$=\dfrac{\exp\left(\sum_{j=0}^{J-1}f_{\bm{\theta}}(T_{c}\bm{x})[Qj+q]\right)}{\sum_{q^{\prime}=0}^{Q-1}\exp\left(\sum_{j=0}^{J-1}f_{\bm{\theta}}(T_{c}\bm{x})[Qj+q^{\prime}]\right)}.$ (1)
For brevity, we do not recall the dependency of $\bm{y}_{\bm{\theta}}$ upon $\bm{x}$ nor $c$. Equation 1 resembles the extraction of chroma features [26], hence the proposed name of ChromaNet.
@@ 3.3 DFT over key signature profiles
With $Q=12$, the discrete Fourier transform (DFT) of the KSP $\bm{y}_{\bm{\theta}}$ is
@@ 3.4 Cross-power spectral density (CPSD)
Let us split the CQT matrix $\bm{x}$ into two disjoint time segments of equal length: $\bm{x}=(\bm{x}_{\mathrm{A}},\bm{x}_{\mathrm{B}})$. We denote the ChromaNet response for A by $\bm{y}_{\bm{\theta},{\mathrm{A}}}=(g\circ f_{\bm{\theta}})(T_{c}\bm{x})$ idem for B. The circular cross-correlation between $\bm{y}_{\bm{\theta},{\mathrm{A}}}$ and $\bm{y}_{\bm{\theta},{\mathrm{B}}}$ is
@@ 4.1 Training on real-world unlabeled data
The self-supervised procedure above learns an approximately equivariant mapping from CQT to key signature profiles (KSP). After training, we observe informally that for each input $\bm{x}$, most of the softmax activation in the KSP $\bm{y}_{\bm{\theta}}$ is concentrated on a single pitch class. In other words, the loss $\mathcal{L}^{\mathrm{CPSD}}$ (Equation 9) is sparsity-promoting.
@@ 4.5 Results on key signature estimation
Feature engineering 1599 981 38%
STONE $(\omega=7)$ 3587 1225 77%
STONE $(\omega=1)$ 3883 920 79%
Supervised SOTA [21] 4090 741 81%
@@ 7 Conclusion
STONE learns key signature profiles (KSP) via equivariant self-supervised learning in the time–frequency domain. We have seen than a semi-supervised extension of STONE (semi-TONE) reduces expert annotation by 90% less at no loss of MIREX score compared to the fully supervised variant (sup-TONE). The primary limitation of our work resides in the inability of the STONE objective (CPSD, i.e., cross-power spectral density) to distinguish major keys from minor keys.re work will study how STONE can be adapted to other pitch-relative MIR tasks.